OpenStudy (anonymous):

a scientist uses a scale Y for measuring temperature. In this scale, pure water melts at 10oY and boils at 130oY at standard atmospheric pressure. he finds that sodium boils at 127oY. Express this temperature in degrees celsius.

OpenStudy (anonymous):

so what am i sopos to do on here

OpenStudy (anonymous):

sketch or eplain pls

OpenStudy (anonymous):

what do do on this question

OpenStudy (unklerhaukus):

In Celsius there are 100 degrees between the melting and boiling of water

OpenStudy (unklerhaukus):

in the Y scale there are 120 degrees so \[120°Y=100°C\] \[1°Y=5/6 °C\]

OpenStudy (unklerhaukus):

\[127°Y=130°Y-3°Y\] \[=100°C-3\times5/6°C\] \[=97.5°C\]

OpenStudy (marco26):

 solving for x: (0-100)/(x-100)=(10-130)/(127-130) x=97.5 degree centigrade